OpenStudy (anonymous):

I'd love someone to demonstrate how to solve this problem step by step. :) (4)/(x²y) - (-5)/(xy²)

OpenStudy (anonymous):

\[\frac{4}{x^2y}-\frac{-5}{xy^2}\] Just like when you add or subtract regular old fractions, you need to have a common denominator. You look at the denominators and figure out what they have in common. They both have an x and a y. That tells you that your common denominator will have an x and a y.

OpenStudy (anonymous):

You also look at what they have that's different. Notice that the first one has an x^2 instead of just an x, and the second one has a y^2 instead of just a y. That tells you that you need another x and another y in your denominator. So your common denominator is going to be \[x^2y^2\] Okay so far?

OpenStudy (anonymous):

Yep. :)

OpenStudy (anonymous):

Okay, so we think, "What do we multiply the bottom by to get x^2y^2?" The only thing the first one is missing is another y. So we are going to multiply the top and the bottom by y. \[\frac{4}{x^2y}\frac{y}{y}=\frac{4y}{x^2y^2}\] Okay to there?

OpenStudy (anonymous):

I think I have it figured out now. Your explanation was really helpful! Thank you. :)

OpenStudy (anonymous):

:) No problem.